OpenStudy (anonymous):

Hillary borrowed the same amount of money from a bank and a friend. The table below shows the amount, in dollars, that Hillary would owe them after different numbers of years: Year 1 2 3 4 Bank 203.50 207 210.50 214 Friend 203 206.05 209.14 212.28 Which statement is true about the money Hillary would owe the bank and her friend after 30 years?

OpenStudy (anonymous):

She would owe the bank more money. She would owe her friend more money She would owe both the same amount of money. She would owe her friend twice the amount she borrowed.

OpenStudy (anonymous):

I just need to know how to find the percentage you multiply 1000 by to get the next year I don't know how to find the x

OpenStudy (anonymous):

because for the bank its constant plus 3.50 but how do I figure out how much money the friend is increasing by. by what decimal

OpenStudy (anonymous):

@SolomonZelman @shamil98

OpenStudy (the_fizicx99):

What's the name of the lesson?

OpenStudy (anonymous):

4.07 exploring linear and exponential growth

OpenStudy (the_fizicx99):

Ok so her friend's loan is exponential, the bank's is linear. An exponential will always exceed a linear function.

OpenStudy (anonymous):

well I know that but how do I figure out the percentage its increasing by so on year 30 I can figure it out

OpenStudy (anonymous):

The friend is increasing the debt by the product (1.015)^n where n is the number of years, as you can tell by taking the ratio of values of adjacent years. This is 1.5% compounded yearly, so in 29 more years after year 1 she will owe (203)(1.015^29= 203(1.54)=312.62 Compare this with the bank where her debt will be 203.50 + 29(3.50). = ?

OpenStudy (anonymous):

bank is 308.5 and friend is 317 for the 30th year thank you so much but my question was howd you figure out the percentage for the friend was increasing 1.015

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

I took the ratio of some adjacent year values and found they were the same, 1.015.